Determine the density of the earth assuming the earth is sphere. Take the radius of the earth r=6.37×10^6

And the mass of the earth m=5.98×10^24


Show your work


1
Expert's answer
2021-02-15T17:38:06-0500

By definition, the density is:


"\rho = \dfrac{m}{V}"

where "m = 5.98\times 10^{24}kg" is the mass of the Earth, and "V" is its volume. Considering the Earth to be a ball with radius "r = 6.67\times 10^6m", the volume is:


"V = \dfrac{4}{3}\pi r^3"

Thus, the density will be:


"\rho = \dfrac{3m}{4\pi r^3}\\
\rho = \dfrac{3\cdot 5.98\times 10^{24}}{4\pi (6.67\times 10^6)^3} \approx 4.81\times 10^3\space \dfrac{kg}{m^3}"

Answer. "4.81\times 10^3\space \dfrac{kg}{m^3}"